Limit x approach infinity x[ ( 1 + a/x )^ 1+ 1/x - x^-1/x ( x+ a ) ]


1
Expert's answer
2021-01-12T14:49:41-0500

"\lim \limits_{x\to \infty} x[ ( 1 + \frac{a}{x} )+ \frac{1}{x} - \frac{x^{-1}}{x} ( x+ a ) ]=""\lim \limits_{x\to \infty} [ ( x + a )+ 1 - ( 1+ \frac{a}{x} ) ]=" "\lim \limits_{x\to \infty} [ x + a - \frac{a}{x} ]=\infty"

As we know "\lim \limits_{x\to \infty} \frac{1}{x}=0" and "\lim \limits_{x\to \infty} x=\infty". so the expression increases indefinitely.